# ICYM2I: The illusion of multimodal informativeness under missingness

### v2

@@ Appendix H Compute infrastructure
All experiments were performed on a server with an AMD EPYC 7313 CPU, 256 GB of memory, and two NVIDIA RTX A6000 GPUs, as well as a server with an Intel Xeon E5-2640 CPU, 128 GB of memory, and a NVIDIA GTX Titan X GPU. Our software stack includes Python 3.12, PyTorch 2.2.1 [95], and standard Python scientific libraries. Chest radiograph embeddings used Tensorflow 2.19 [96] and Tensorflow-Text 2.19 based on the requirements for the ELIXR models [72]. Electrocardiogram embeddings were generated using an environment with Python 3.9 and fairseq-signals 1.0 to match the requirements for fairseq-signals and ECG-FM [71]. Generating embeddings for our structural heart disease data took approximately 10 hours on our server with a Titan X GPU. All synthetic experiments require 12 hours of compute time using one NVIDIA RTX A6000 GPU.

### v3

@@ Appendix H Compute infrastructure
All experiments were performed on a server with an AMD EPYC 7313 CPU, 256 GB of memory, and two NVIDIA RTX A6000 GPUs, as well as a server with an Intel Xeon E5-2640 CPU, 128 GB of memory, and a NVIDIA GTX Titan X GPU. Our software stack includes Python 3.12, PyTorch 2.2.1 (Paszke et al., 2019), and standard Python scientific libraries. Chest radiograph embeddings used Tensorflow 2.19 (Abadi et al., 2015) and Tensorflow-Text 2.19 based on the requirements for the ELIXR models (Xu et al., 2023). Electrocardiogram embeddings were generated using an environment with Python 3.9 and fairseq-signals 1.0 to match the requirements for fairseq-signals and ECG-FM (McKeen et al., 2024). Generating embeddings for our structural heart disease data took approximately 10 hours on our server with a Titan X GPU. All synthetic experiments require 12 hours of compute time using one NVIDIA RTX A6000 GPU.